# Understanding the Cluster LP for Correlation Clustering

### v2

@@ 5.4 $+$\!+\!-+ + - Triangles
Fix budget functions $b^{+}\equiv b^{+}_{\alpha}$ and $b^{-}\equiv b^{-}_{\alpha}$, where $\alpha=\frac{3}{2}$. For any ++\!-\!-+ - - triangle $T$, we have $\mathrm{cost}(T)\leq\Delta(T)$.
@@
Again, we distinguish cases depending on the number of short $+$edges.
@@
For a $+$\!+\!-+ + - triangle $T=(u,v,w)$, assume that $uv$ is a --edge and $uw$, $vw$ are short $+$edges.
$\mathrm{cost}(T)$ $=2\cdot x_{uv}+1,$ (16)
$\Delta(T)$ $=b^{-}(x_{uv})+b^{+}(x_{uw})+b^{+}(x_{vw})\geq b^{-}(x_{uv})+2\cdot b^{+}\left(\frac{x_{uw}+x_{vw}}{2}\right)\geq b^{-}(x_{uv})+2\cdot b^{+}\left({\frac{x_{uv}}{2}}\right).$ (17)
As before, we used the convexity of $b^{+}$ (Lemma 24) together with the triangle inequality, $x_{uv}\leq x_{uw}+x_{vw}$. The derivative of (17) w.r.t. $x_{uv}$ is less than 00 for $x_{uv}\geq 0$,
$-\frac{6x_{uv}(4+3x_{uv})}{(1+x_{uv})^{2}(2+x_{uv})^{2}}\leq 0.$
Hence, (17) is decreasing with $x_{uv}$. Observe that $x_{uv}\leq x_{uw}+x_{vw}\leq\frac{2}{3}$. Since the cost is increasing with $x_{uv}$ we only need to check the case $x_{uv}=\frac{2}{3}$,
$\Delta(T)-\mathrm{cost}(T)\geq b^{-}\left(\frac{2}{3}\right)+2\cdot b^{+}\left({\frac{1}{3}}\right)-\frac{7}{3}=\frac{1}{15}.$
@@
For a $+$\!+\!-+ + - triangle $T=(u,v,w)$, assume that $uv$ is a --edge, $uw$ is a short $+$edge and $vw$ is a long $+$edge.
$\mathrm{cost}(T)$ $=x_{uv}+y_{vw}+y_{vw}+y_{uv}-2y_{vw}y_{uv}=1+2y_{wv}(1-y_{uv})=1+2(1-x_{wv})x_{uv},$
$\Delta(T)$ $=b^{+}(x_{vw})+b^{-}(x_{uv})+b^{+}(x_{uw})(y_{vw}+y_{uv}-y_{vw}y_{uv})$
$=b^{+}(x_{vw})+b^{-}(x_{uv})+b^{+}(x_{uw})(2-x_{vw}-x_{uv}-(1-x_{vw})(1-x_{uv})).$
@@
$x_{vw}\geq x_{uv}$.
$\Delta(T)$ $\geq b^{+}(x_{uv})+b^{-}(x_{uv})=3.$
$\mathrm{cost}(T)$ is always less than $3$.
@@
$x_{vw}\leq x_{uv}$. We have $x_{uw}\geq x_{uv}-x_{vw}$ by triangle inequality. Since $b^{+}(x)$ is increasing by Lemma 24,
$\Delta(T)\geq b^{+}(x_{vw})+b^{-}(x_{uv})+b^{+}(x_{uv}-x_{vw})(2-x_{vw}-x_{uv}-(1-x_{vw})(1-x_{uv})).$
For a fixed $z\coloneqq x_{uv}$, define $x\coloneqq x_{vw}\geq\frac{1}{3}$. We need to argue that the difference
$\Delta(T)-\mathrm{cost}(T)=b^{+}(x)+b^{-}(z)+b^{+}(z-x)(2-x-z-(1-x)(1-z))-1-2(1-x)z$ (18)
is non-negative.
@@
The derivative of (18) w.r.t. $x$ is
$\frac{2(6x^{5}z+x^{4}(4-15z)z+4x^{3}z(-3-5z+3z^{2})+x^{2}z(-11+7z+22z^{2}-3z^{3}))}{(1+x)^{2}(1-x+z)^{2}}$
$+\frac{2(-z(5+z+2z^{2}+3z^{3})+x(12+6z+14z^{2}+8z^{3}-6z^{4}))}{(1+x)^{2}(1-x+z)^{2}},$ (19)
which is non-negative as long as the numerator is non-negative. The numerator is a concave function in $x$. Indeed, the second derivative of the numerator in (19) w.r.t. $x$ is
$4z(-11+60x^{3}+x^{2}(24-90z)+7z+22z^{2}-3z^{3}+12x(-3-5z+3z^{2})).$ (20)
We want to show that (20) is non-positive. This is equivalent to
$-11+60x^{3}+x^{2}(24-90z)+7z+22z^{2}-3z^{3}+12x(-3-5z+3z^{2})\leq 0.$ (21)
Note that (21) is a convex function in $x$. Indeed, the second derivative w.r.t. $x$ of (21) is equal to
$12(4+30x-15z)\geq 0$
The inequality holds true since either $z\geq 2/3$ and $x\geq z-\frac{1}{3}$ or $z\leq 2/3$. Hence, (21) is maximized for either $x=z$, $x=1/3$ or $x=z-1/3$.
$x=z,$ (21)$=-11-29z-14z^{2}+3z^{3}\leq 0,$
$x=\frac{1}{3},$ (21)$=-\frac{163}{9}-23z+34z^{2}-3z^{3}\leq 0,$
$x=z-\frac{1}{3},$ (21)$=\frac{13}{9}-15z-26z^{2}+3z^{3}\leq 0.$
All inequalities above hold for $\frac{1}{3}\leq z\leq 1$. Thus, the numerator of (19) is concave and minimized for either $x=1$, $x=1/3$ or $x=z-1/3$.
$x=z,$ $\quad 2z(7+5z+z^{2})\geq 0,$
$x=\frac{1}{3},$ $\quad\frac{2}{27}(108-124z+95z^{2}+96z^{3}-144z^{4})\geq 0,$
$x=z-\frac{1}{3},$ $\quad-8+\frac{688}{81}z+\frac{86}{9}z^{2}+\frac{22}{3}z^{3}+\frac{14}{3}z^{4}\geq 0.$
The first two inequalities hold for $\frac{1}{3}\leq z\leq 1$. The last inequality holds for $\frac{2}{3}\leq z\leq 1$. We can conclude that (18) is indeed increasing with $x$. $\Diamond$
Therefore, we can assume that $x=1/3$ if $z\leq 2/3$ or $x=z-1/3$.
Case 2.1 $x=1/3.$
(18)$=\frac{1}{2}+b^{-}(z)+b^{+}\left(z-\frac{1}{3}\right)\left(1-\frac{1}{3}z\right)-1-\frac{4}{3}z.$ (22)
The derivative of (22) is equal to
$\frac{-2(-22+124z+494z^{2}+510z^{3}+153z^{4})}{3(1+z)^{2}(2+3z)^{2}}.$
For $z\geq x=1/3$, we have that,
$-22+124z+494z^{2}+510z^{3}+153z^{4}\geq 95.$
Hence, (22) is decreasing with $z$. For $z=2/3$, (22)=2/5italic-(22italic-)$=2/5$.
Case 2.2 $x=z-1/3.$
(18)$=b^{+}\left(z-\frac{1}{3}\right)+b^{-}(z)+\frac{1}{2}\left(\frac{7}{3}-2z-\left(\frac{4}{3}-z\right)(1-z)\right)-1-2\left(\frac{4}{3}-z\right)z$ (23)
The derivative of (23) is equal to
$\frac{-80-148z-65z^{2}+72z^{3}+135z^{4}+54z^{5}}{2(1+z)^{2}(2+3z)^{2}}.$
For $0\leq z\leq 1$, we have that,
$-80-148z-65z^{2}+72z^{3}+135z^{4}+54z^{5}\leq 0.$
Hence, (23) is decreasing with $z$. For $z=1$, (23)=110italic-(23italic-)$=\frac{1}{10}$.
@@
For a $+$\!+\!-+ + - triangle $T=(u,v,w)$, assume that $uv$ is a --edge and $uw$, $vw$ are long $+$edges. We will proceed similar to [CLLN23]. For --edges we use a coefficient of $3$, $b^{-}(x)\geq 3(1-x)$. For $+$edges define the coefficient $f(x)=\frac{6x}{1+x}$, $b^{+}(x)=f(x)\cdot x$.
$\Delta(T)$ $\geq 3(y_{uw}+y_{vw}-y_{uvw})y_{uv}+f(x_{vw})(y_{uw}+y_{uv}-y_{uw}y_{uv})(1-y_{vw})$
$\quad+f(x_{uw})(y_{vw}+y_{uv}-y_{vw}y_{uv})(1-y_{uw})$
$=f(x_{vw})y_{uw}+f(x_{uw})y_{vw}+(f(x_{uw})+f(x_{a,c}))y_{uv}+(3-f(x_{uw})-f(x_{vw}))(y_{uw}+y_{vw})y_{uv}$
$\quad-(f(x_{uw})+f(x_{vw}))y_{uw}y_{vw}(1-y_{uv})-3y_{uvw}y_{uv}.$
$\mathrm{cost}(T)$ $=y_{uvw}+y_{uw}+y_{uv}-2y_{uw}y_{uv}+y_{vw}+y_{uv}-2y_{vw}y_{uv}=y_{uvw}+y_{uw}+y_{vw}+2y_{uv}-2(y_{uw}+y_{vw})y_{uv}.$
The difference between the cost and the budget is at least,
$\Delta(T)-\mathrm{cost}(T)\geq(f(x_{vw})-1)y_{uw}+(f(x_{uw})-1)y_{vw}+(f(x_{uw})+f(x_{a,c})-2)y_{uv}$
$+(5-f(x_{uw})-f(x_{vw}))(y_{uw}+y_{vw})y_{uv}-(f(x_{uw})+f(x_{vw}))y_{uw}y_{vw}(1-y_{uv})-3y_{uvw}y_{uv}-y_{uvw}.$ (24)
We can decrease $y_{uv}$ until $y_{uv}=y_{uvw}$.
@@
We will show that the derivative of (24) w.r.t. $y_{uv}$ is non-negative. The derivative of (24) w.r.t. $y_{uv}$ is equal to
$(5-f(x_{uw})-f(x_{vw})))(y_{uw}+y_{vw})+(f(x_{uw})+f(x_{vw})-2)+(f(x_{uw})+f(x_{vw}))(y_{uw}y_{vw})-3y_{uvw}.$ (25)
Observe that $f(x)\leq 3$ and $f(x)\geq 2$ if $x\geq\frac{1}{2}$. Hence, if $x_{uw},x_{vw}\geq\frac{1}{2}$,
(25)$\geq-(y_{uw}+y_{vw})+2+4(y_{uw}y_{vw})-3y_{uvw}\geq 1+4y^{2}_{uvw}-3y_{uvw}\geq 0.$
Note that $f(x)\leq 2$ if $x\leq\frac{1}{2}$ and $f(x)\geq\frac{3}{2}$ if $x\geq\frac{1}{3}$. Hence, If $\frac{1}{3}\leq x_{uw}\leq\frac{1}{2}$ and $x_{vw}\geq\frac{1}{3}$, then,
(25)$\geq 1+3(y_{uw}y_{vw})-3y_{uvw}\geq 1+3y^{2}_{uvw}-3y_{uvw}+1+3y_{uvw}(1-y_{uvw})\geq 0.$
@@
Substituting $y_{uvw}=y_{uv}$ into (24) yields,
$(f(x_{vw})-1)y_{uw}+(f(x_{uw})-1)y_{vw}+(5-f(x_{uw})-f(x_{vw}))(y_{uw}+y_{vw})y_{uv}$
$+(f(x_{uw})+f(x_{vw})-3)y_{uv}-(f(x_{uw})+f(x_{vw}))y_{uw}y_{vw}(1-y_{uv})-3y^{2}_{uv}.$ (26)
We have to show that (26) is non-negative. For $y_{uv}$ we have the constraints $y_{uv}=y_{uvw}\leq\min\{y_{uw},y_{vw}\}$ and $y_{uv}\geq(y_{uw}+y_{vw}-1)_{+}$. (26) is a quadratic in $y_{uv}$ with leading coefficient $-3$. Hence, (26) is minimized either by $y_{uv}=(y_{uw}+y_{vw}-1)_{+}$ or by $y_{uv}=\min\{y_{uw},y_{vw}\}$.
Case 1: $y_{uv}=0$.
In this case,
(26)$=(f(x_{vw})-1)y_{uw}+(f(x_{uw})-1)y_{vw}-(f(x_{uw})+f(x_{vw}))y_{uw}y_{vw}.$ (27)
If $x_{uw},x_{vw}\geq\frac{1}{2}$, then,
@@
$\geq 2(y_{uw}-y_{uw}y_{vw})+2(y_{vw}-y_{uw}y_{vw})-(y_{uw}+y_{vw})$
$=y_{uw}+y_{vw}-4y_{uw}y_{vw}$
$\geq 0.$
For the last inequality we used that $y_{uw},y_{vw}\leq\frac{1}{2}$ and hence, $y_{uw}\geq 2y_{uw}y_{vw}$. Otherwise, we can assume w.l.o.g. that $x\coloneqq x_{vw}\leq\frac{1}{2}\leq 1-x\leq x_{uw}=:y$, since $1\geq y_{uv}+y_{uw}+y_{vw}-2y_{uvw}=y_{uv}+y_{uw}$.
@@
The derivative of (27) w.r.t. $y$ is
$\frac{(2y+y^{2})(7+x-12x^{2})+6x^{2}-x-1}{(y+1)^{2}(x+1)}\geq 0.$
The inequality holds since for $y\geq\frac{1}{2}$ and $x\leq\frac{1}{2}$, the numerator is bounded by
$(2y+y^{2})(7+x-12z^{2})+6x^{2}-x-1\geq\frac{31}{4}+\frac{1}{4}x-9x^{2}\geq 0.$

### v3

@@ 4.3 Using Independently Sampled Clusters to Construct Cluster LP Solution
Fix budget functions $\mathrm{budget}^{+}\equiv\mathrm{budget}^{+}_{\alpha}$ and $\mathrm{budget}^{-}\equiv\mathrm{budget}^{-}_{\alpha}$, where $\alpha=1.56$. For any +++\!+\!- triangle $T$, we have $\mathrm{cost}(T)\leq\Delta(T)$.
@@
Again, we distinguish cases depending on the number of short ++edges.
@@
For a +++\!+\!- triangle $T=(u,v,w)$, assume that $uv$ is a -edge and $uw$, vwvw are short ++edges. As usual, let $x_{uv}=a,x_{uw}=b$ and $x_{vw}=c$.
$\mathrm{cost}(T)$ $=2\cdot a+1,$
$\Delta(T)$ $=\mathrm{budget}^{-}(a)+\mathrm{budget}^{+}(b)+\mathrm{budget}^{+}(c)\geq\mathrm{budget}^{-}(a)+2\cdot\mathrm{budget}^{+}\left(\frac{b+c}{2}\right)$
$\geq\mathrm{budget}^{-}(a)+2\cdot\mathrm{budget}^{+}\left({\frac{a}{2}}\right)~=\frac{2\alpha}{2-\alpha}\left(\frac{(1+2a)(1-a)}{2(1+a)}+\frac{a^{2}}{2+a}\right).$ (23)
As before, we used the convexity of $\mathrm{budget}^{+}$ (Lemma 33) together with the triangle inequality, $a\leq b+c$. The derivative of (23) w.r.t. $a$ is less than 0 for $a\geq 0$,
$\frac{2\alpha}{2-\alpha}\left(\frac{a(3a+4)}{(1+a)^{2}(2+a)^{2}}\right)\leq 0.$
Hence, (23) is decreasing with $a$. Observe that $a\leq b+c\leq 2\tau$. Since the cost is increasing with $a$ we only need to check the case $a=2\tau$,
$\Delta(T)-\mathrm{cost}(T)$ $\geq\mathrm{budget}^{-}\left(2\tau\right)+2\cdot\mathrm{budget}^{+}\left(\tau\right)-4\tau-1$
$\geq\mathrm{budget}^{-}\left(2\tau\right)+4\tau-4\tau-1\geq.02.$
Notice that we use Claim 34 to lower bound $\mathrm{budget}^{+}(\tau)$.
@@
For a +++\!+\!- triangle $T=(u,v,w)$, assume that $uv$ is a -edge, uwuw is a short ++edge and $vw$ is a long ++edge.
$\mathrm{cost}(T)$ $=a+y_{vw}+y_{vw}+y_{uv}-2y_{vw}y_{uv}=1+2y_{vw}(1-y_{uv})=1+2(1-x_{vw})a=1+2(1-c)a.$
$\Delta(T)$ $=\mathrm{budget}^{+}(c)+\mathrm{budget}^{-}(a)+\mathrm{budget}^{+}(b)(y_{vw}+y_{uv}-y_{vw}y_{uv})$
$=\mathrm{budget}^{+}(c)+\mathrm{budget}^{-}(a)+\mathrm{budget}^{+}(b)(2-c-a-(1-c)(1-a))$
$=\mathrm{budget}^{+}(c)+\mathrm{budget}^{-}(a)+\mathrm{budget}^{+}(b)(1-ac).$
@@
$c\geq a$.
$\Delta(T)$ $\geq\mathrm{budget}^{+}(a)+\mathrm{budget}^{-}(a)=\frac{\alpha}{2-\alpha}\geq 3,$
while $\mathrm{cost}(T)$ is always less than $3$.
@@
$c\leq a$. We have $b\geq a-c$ by triangle inequality. Since $b\leq\tau$, we also have $a-c\leq\tau$. Since $\mathrm{budget}^{+}(x)$ is increasing by Lemma 33,
$\Delta(T)\geq\mathrm{budget}^{+}(c)+\mathrm{budget}^{-}(a)+\mathrm{budget}^{+}(a-c)(1-ac).$
Define $C_{\alpha}=\frac{\alpha}{2-\alpha}$. For a fixed $x\coloneqq c\geq\tau$, define $z\coloneqq a$. For -edges we can use a coefficient of $C_{\alpha}$, since $\mathrm{budget}^{-}(x)\geq C_{\alpha}(1-x)$. We need to argue that the difference on line (24) is non-negative.
$\Delta(T)-\mathrm{cost}(T)$ $\geq\mathrm{budget}^{+}(x)+\mathrm{budget}^{-}(z)+\mathrm{budget}^{+}(z-x)(1-xz)-1-2(1-x)z$ (24)
$\geq\mathrm{budget}^{+}(x)+C_{\alpha}(1-z)+\mathrm{budget}^{+}(z-x)(1-xz)-1-2(1-x)z$ (25)
$=2C_{\alpha}\frac{x^{2}}{1+x}+C_{\alpha}(1-z)+2C_{\alpha}\frac{(z-x)^{2}}{1+z-x}(1-xz)-1-2(1-x)z.$ (26)
Define
$f(z):=C_{\alpha}(1-z)+2C_{\alpha}\frac{(z-x)^{2}}{1+z-x}(1-xz)-1-2(1-x)z.$
@@
We will show that the derivative of $f$ is negative. We have
$f^{\prime}(z)$ $=C_{\alpha}\left(2x^{2}-4xz+2x+1+\frac{2(x-1)x-2}{(1+z-x)^{2}}\right)+2(x-1)$
$\leq C_{\alpha}\left(-2xz+2x+1+\frac{2(x-1)x-2}{(1+z-x)^{2}}\right).$
We will show that $g(z)\coloneqq\left(-2xz+2x+1+\frac{2(x-1)x-2}{(1+z-x)^{2}}\right)<0$ and thus, $f^{\prime}(z)<0$. We have,
$g^{\prime}(z)=-\frac{4(x^{2}-x-1)}{(1+z-x)^{3}}-2x,$
and,
$g^{\prime\prime}(z)=\frac{12(x^{2}-x-1)}{(1+z-x)^{4}}.$
Note that $g^{\prime\prime}(z)\leq 0$ since $C>0$, the numerator is negative for $x\in[0,1]$ and the denominator is always positive. Thus, $g^{\prime}(z)$ is decreasing and minimized for $z=1$,
$g^{\prime}(1)=-\frac{4(x^{2}-x-1)}{(2-x)^{3}}-2x\geq 0,$
for all $x\in[0,1]$. Hence, $g(z)$ is increasing and maximized for $z=\min\{1,x+\tau\}$. We have,
$g(1)=\left(-2x+2x+1+\frac{2(x-1)x-2}{(2-x)^{2}}\right)<0$
for $x\geq 1-\tau$ and,
$g(x+\tau)=\left(-2x(x+\tau)+2x+1+\frac{2(x-1)x-2}{(1+\tau)^{2}}\right)<0,$
for $x\in[0,1]$. $\Diamond$
Now we return to our goal, which is to show that
$2C_{\alpha}\frac{x^{2}}{1+x}+C_{\alpha}(1-z)+2C_{\alpha}\frac{(z-x)^{2}}{1+z-x}(1-xz)-1-2(1-x)z\geq 0.$ (27)
By Claim 40, we can assume that $z=\min\{1,\tau+x\}$. First, we consider the case in which $\tau+x\geq 1$, in which case we assume that $z=1$. Thus, we have
[…]
for $x\geq 1-\tau$.
Next, we consider the case in which $z=x+\tau<1$,
[…]
for $x\in[0,1]$.
@@
For a +++\!+\!- triangle $T=(u,v,w)$, assume that $uv$ is a -edge and $uw$, vwvw are long ++edges. Define $C_{\alpha}=\frac{\alpha}{2-\alpha}$. For -edges we can use a coefficient of $C_{\alpha}$, since $\mathrm{budget}^{-}(x)\geq C_{\alpha}(1-x)$. For ++edges define $f(x)=\frac{2C_{\alpha}\cdot x^{2}}{1+x}$ so that $\mathrm{budget}^{+}(x)=f(x)$. Recall that $a=x_{uv},b=x_{uw}$ and $c=x_{vw}$. Without loss of generality, we assume that $b\leq c$, which implies $y_{uw}\geq y_{vw}$.
$\Delta(T)$ $=C_{\alpha}(y_{uw}+y_{vw}-y_{uvw})y_{uv}+f(c)(y_{uw}+y_{uv}-y_{uw}y_{uv})+f(b)(y_{vw}+y_{uv}-y_{vw}y_{uv})$
$\geq C_{\alpha}(y_{uw}+y_{vw}-y_{uvw})y_{uv}+2c(y_{uw}+y_{uv}-y_{uw}y_{uv})+2b(y_{vw}+y_{uv}-y_{vw}y_{uv}).$
$\mathrm{cost}(T)$ $=y_{uvw}+y_{uw}+y_{uv}-2y_{uw}y_{uv}+y_{vw}+y_{uv}-2y_{vw}y_{uv}$
$=y_{uvw}+y_{uw}+y_{vw}+2y_{uv}-2(y_{uw}+y_{vw})y_{uv}.$
The inequality follows from Claim 34. Using $y_{uv}=1-a,y_{uw}=1-b,y_{vw}=1-c$ and $y=y_{uvw}$, we have
$\Delta(T)$ $\geq C_{\alpha}(y_{uw}+y_{vw}-y_{uvw})y_{uv}+2c(y_{uw}+y_{uv}-y_{uw}y_{uv})+2b(y_{vw}+y_{uv}-y_{vw}y_{uv})$
$=C_{\alpha}(2-b-c-y)(1-a)+2c+2b-4abc.$
For the difference we get,
$\Delta(T)-\mathrm{cost}(T)\geq C_{\alpha}(2-b-c-y)(1-a)+3c+3b+2a-4abc-y-4+2(2-b-c)(1-a)$
$=2C_{\alpha}+(1-C_{\alpha})b+(1-C_{\alpha})c+(-C_{\alpha}-1+C_{\alpha}a)y+(-2C_{\alpha}-2)a+(C_{\alpha}+2)ab+(C_{\alpha}+2)ac-4abc.$
Define the previous line to be the function $g(a,b,c,y)$. Our goal is to that that
$\Delta(T)-\mathrm{cost}(T)\geq g(a,b,c,y)\geq 0.$
To minimize $g$, we can set $y$ as large as possible, since $y$ has a negative coefficient in $g$. So we have
$y=\min\{y_{uv},y_{vw}\}=\min\{1-a,1-c\},$
because we assumed $b\leq c$ and hence $y_{vw}\leq y_{uw}$.
Case 1: $a\leq b\leq c$.
First, we consider the case in which $a\leq b\leq c$. In this case, we can set y=1cy=1-c. So we have
$g_{1}(a,b,c)$ $:=2C_{\alpha}+(1-C_{\alpha})(b+c)+(-C_{\alpha}-1+C_{\alpha}a)(1-c)+(-2C_{\alpha}-2)a+(C_{\alpha}+2)ab+(C_{\alpha}+2)ac-4abc$
$=(1-C_{\alpha})b+(C_{\alpha}-1)+(2-C_{\alpha}a)c+(-C_{\alpha}-2)a+(C_{\alpha}+2)ab+(C_{\alpha}+2)ac-4abc$
$=(1-C_{\alpha})b+(C_{\alpha}-1)+2c+(-C_{\alpha}-2)a+(C_{\alpha}+2)ab+2ac-4abc.$
@@
The derivative of $g_{1}$ with respect to $c$ is
$2+2a-4ab\geq 2+2a-4a=2(1-a)\geq 0.$
Here, we used that $b\leq 1$. $\Diamond$
So we can make $c$ as small as possible and we thus set $c=b$.
$g_{1}(a,b,c)\geq g_{1}(a,b,b)=(3-C_{\alpha})b+(C_{\alpha}-1)+(-C_{\alpha}-2)a+(C_{\alpha}+4)ab-4ab^{2}.$
Claim 42.
$g_{1}(a,b,b)$ is decreasing with $a$.
Proof..
The derivative is
$C_{\alpha}-2+C_{\alpha}b+4b-4b^{2}.$
The above quadratic is maximized for $b=\frac{C_{\alpha}+4}{8}$,
$C_{\alpha}-2+C_{\alpha}b+4b-4b^{2}\leq\frac{1}{16}(C_{\alpha}^{2}-8C_{\alpha}-16)\leq 0$
@@
Now we can increase $a$ as much as possible and set $a=b$,
$g_{1}(b,b,b)=(C_{\alpha}-1)+(1-2C_{\alpha})b+(C_{\alpha}+4)b^{2}-4b^{3}.$ (30)
Claim 43.
$g_{1}(b,b,b)$ is decreasing in $b$.
Proof..
The derivative is,
$1-2C_{\alpha}+(2C_{\alpha}+8)b-12b^{2}.$
The above quadratic is maximized for $b=\frac{C+4}{12}$,
$1-2C_{\alpha}+(2C_{\alpha}+8)b-12b^{2}\leq\frac{1}{12}(C_{\alpha}^{2}-16C_{\alpha}+28)\leq 0$
. $\Diamond$
Thus, we have,
$g_{1}(b,b,b)\geq g_{1}(1,1,1)=0$
Case 2: $b\leq a\leq c$.
Our goal is to show that in this case, $g_{1}(a,b,c)\geq 0$. Claim 41 still applies. When we decrease $c$ as much as possible, we have $c=a$. So we have
$g_{1}(a,b,c)\geq g_{1}(a,b,a)=g_{2}(a,b)$ $:=(1-C_{\alpha})b+(C_{\alpha}-1)-C_{\alpha}a+(C_{\alpha}+2)ab+2a^{2}-4a^{2}b.$
Claim 44.
$g_{2}$ is decreasing in $b$.
Proof..
The derivative is
$1-C_{\alpha}+(C_{\alpha}+2)a-4a^{2}.$
The above quadratic is maximized for $a=\frac{C_{\alpha}+2}{8}$,
$1-C_{\alpha}+(C_{\alpha}+2)a-4a^{2}\leq\frac{1}{16}(C_{\alpha}^{2}-12C_{\alpha}+20)\leq 0.$
$\Diamond$
So we increase $b$ as much as possible, setting $b=a$. Then we have
$g_{2}(a,b)\geq g_{2}(a,a)=g_{3}(a)$ $:=a+(C_{\alpha}-1)-2C_{\alpha}a+(C_{\alpha}+4)a^{2}-4a^{3}.$
Claim 45.
$g_{3}$ is decreasing in $a$.
Proof..
The derivative is
$1-2C_{\alpha}+2(4+C_{\alpha})a-12a^{2}.$
The above quadratic is maximized for $a=\frac{C_{\alpha}+4}{12}$,
$1-2C_{\alpha}+2(4+C_{\alpha})a-12a^{2}\leq\frac{1}{12}(C_{\alpha}^{2}-16C_{\alpha}+28)\leq 0.$
$\Diamond$
By Claim 45, we have $g_{3}(a)\geq g_{3}(1)$, and we have that $g_{3}(1)=0$.
Case 3: $b\leq c\leq a$.
Recall that our goal is to prove the following quantity is $\geq 0$.
$2C_{\alpha}+(1-C_{\alpha})b+(1-C_{\alpha})c+(-C_{\alpha}-1+C_{\alpha}a)y+(-2C_{\alpha}-2)a+(C_{\alpha}+2)ab+(C_{\alpha}+2)ac-4abc.$
In this case, we can set y=1ay=1-a, so we have
$C_{\alpha}-1+(1-C_{\alpha})b+(1-C_{\alpha})c-a-C_{\alpha}a^{2}+(C_{\alpha}+2)ab+(C_{\alpha}+2)ac-4abc.$ (31)
Claim 46.
(31) is decreasing with $a$.
Proof..
The derivative w.r.t. $a$ is
$-1-2C_{\alpha}a+(C_{\alpha}+2)(b+c)-4bc.$ (32)
The coefficient of $a$ in 32 is $2C_{\alpha}\leq 0$. Similar, the coefficient of $b$ in 32 is $C+2-4c\geq C-2\geq 0$. Thus, we can assume that $a=b=c$,
@@
$\Diamond$
Thus, to minimize (31), we can make $a$ as large as possible, setting $a=\min\{b+c,1\}$. Let us consider the first subcase in which $b+c>1$ and so $a=1$. In this case, (31) becomes
$-2+3(b+c)-4bc=-(2b-1)(2c-1)+b+c-1\geq-(2b-1)(2c-1).$ (33)
Note that $c\geq\frac{1}{2}$ since $b+c\geq 1$ and $c\leq b$. Thus 33 is decreasing with $b$,
$(2b-1)(2c-1)\geq-(2(1-c)-1)(2c-1)=(1-2c)^{2}\geq 0$
The second subcase is when $b+c\leq 1$ and we set $a=b+c$ or equivalently, c=abc=a-b. Then (31) becomes
$C_{\alpha}-1+(4b^{2}-C_{\alpha})a+(2-4b)a^{2}.$ (34)
@@
The derivative of (34) w.r.t. $a$ is,
$4b^{2}-C_{\alpha}+2(2-4b)a.$
Note that $b\leq\frac{1}{2}$ since $b+c\leq 1$ and $b\leq c$. Thus the derivative is increasing with $a$ and maximized for $a=1$,
$4b^{2}-C_{\alpha}+2(2-4b)a\leq 4b^{2}-8b-C_{\alpha}+4.$
which is decreasing for $b\leq 1$ and thus maximized for $b=\tau$,
$4\tau^{2}-8\tau-C_{\alpha}+4\leq 0$